# Understanding and predicting adsorption energetics on monolayer transition metal dichalcogenides

### v1

@@ III.2 Rationalizing trends in adsorption energy
To rationalize the observed trends, we employ the d-band center model, which combines the effects of electron filling and the formation of bonding and anti-bonding orbitals between the adsorbate and the substrate (also see Section II.2. Physically, the strength of interaction between the adsorbate and substrate is related to the filling of the anti-bonding states between the re-normalized valence states of the substrate and adsorbate. Because anti-bonding states generally lie energetically above the electronic band edge, the center of the d-states is an indicator of the bond strength between adsorbate and substrate. The center of the d-band is often referenced to the Fermi level, which determines the energy level of half occupation and thus delineates the filled and unfilled electronic states. Thus, a higher in energy d-band center relative to the Fermi level is indicative of a stronger adsorption interaction, as fewer anti-bonding states are occupied. Hence, the d-band model provides a basis for explaining differences and trends in adsorption energetics. Traditionally, this model focuses on the position of the d-band center of the transition metal substrate relative to the Fermi level as the typical adsorbates are organic chemical intermediates involved in catalytic reactions. By contrast, the adsorbates in this study are largely transition metal adatoms on select substrates (MoS2, MoSe2, WS2, WSe2) that show minimal difference in their structural relaxations (Figure 2). In this study, we apply the d-band center model by focusing on the d-band centers of the transition metal adsorbates instead of the TMD substrates.
Figure 4: Comparison of the d-band centers of the adsorbates, referenced to the Fermi energy, with their respective adsorption energies on MoS2. A d-band center of zero indicates alignment with the Fermi energy. The plots are categorized by specific d-block elements ((a) 3d, (b) 4d, (c) 4d, (d) all together) and color-coded by periodic groups. Linear fits with accompanying R2-values are provided as a guide the eye.
@@
As observed from Figure 4 for MoS2, the early transition metals, such as Hf, Zr, and Ti, have d-band centers shifted closest to the Fermi level when compared to other transition metals. A similar observation can be made for MoSe2, WS2, and WSe2 (see Figures S6, S7, and S8 in the SI). According to the d-band model, this shift of the d-bands closer to the Fermi level correctly anticipates a stronger adsorption energy, as the anti-bonding states would be minimally occupied. In contrast, the late transition metals, such as Cd, Hg, Zn, Au, Cu, and Ag, have d-band centers shifted farther from the Fermi level, indicating lower energy levels for the anti-bonding orbitals, leading to weaker adsorption energy. Consequently, the d-band centers for the re-normalized electronic states of the middle transition metals are shifted by an intermediate amount, resulting in intermediate adsorption energies and strengths. By comparing the shifts of the d-band centers of different adsorbates on the surfaces of TMDs, we can directly correlate the electronic structure of the adsorbates with their adsorption energies. This approach provides a clear understanding of the trends observed across different transition metal adsorbates, highlighting the utility of the d-band center model in rationalizing trends in adsorption properties in resistive switching contexts [75].
Interestingly, we observe that by separating the transition metal adsorbates into 3d, 4d, and 5d metals, there is a more robust correlation between the computed adsorption energy and the d-band center (Figure 4). We hypothesize this to be attributed to the different (increasing) spatial extents of the 3d, 4d, and 5d transition metals, which play a role in the extent of renormalization (i.e., hybridization) between the adsorbate and surface. For example, the 3d transition metals show the lowest correlation between adsorption energy and the d-band center. In contrast, the 4d and 5d orbitals are more spatially extended and exhibit a more linear relationship between the d-band center and adsorption energy, with $R^{2}$ values of 0.78 and 0.95, respectively, for MoS2 (see also Figures S6, S7, and S8 in the SI). For the 3d orbitals, the relationship is less linear, with an $R^{2}$ value of 0.69.
@@
We next turn to crystal orbital Hamilton population (COHP) analysis. COHP analysis provides insights into the degree of bonding and anti-bonding character in the interactions between adsorbates and substrates, helping to further understand the nature of adsorption. By decomposing the electronic density of states (DOS) into its anti-bonding and bonding contributions as a function of energy we obtain the same overall trend as seen using the d-band center model. Using the plots shown in Figure 5, the percentage contribution of anti-bonding interactions can be determined by integrating the Crystal Orbital Hamiltonian Population (COHP) curve for energies below the Fermi level.
Figure 5: Crystal Orbital Hamiltonian Population (COHP) plots of representative transition metal adsorbates from groups 11 (Au), 9 (Co), 6 (Cr), 4 (Zr) on MoS2 are presented. The blue plots show the interactions between the transition metal adsorbate and the molybdenum adsorption site in the monolayer, while the orange plots show the interactions between the transition metal adsorbate and the surrounding sulfur in the monolayer. An estimate of the percent anti-bonding character up to the Fermi level is computed for each adsorbate and corresponds to anti-bonding interactions between the adsorbate and the MoS2 atoms. The highlighted region for AuS is used to illustrate the portion of the COHP curves used to compute the percent anti-bonding character. A similar method is used for each of the different adsorbates. The gold, cobalt, chromium, and zirconium adsorbates exhibit 74.18%, 45.87%, 29.17%, and 9.62% anti-bonding character with sulfur (orange), respectively.
One may naively expect that the adsorption interactions between a TMD and a transition-metal adsorbate would be primarily governed by metal-metal interactions (for example, Mo-Au interaction). To test this whether this is true, we used our COHP analysis to decompose the overall bonding and anti-bonding interactions by species. The COHP analysis shows that Hf, Zr, and Ti have the lowest anti-bonding interactions with their nearest neighbor sulfur atoms (with percent anti-bonding contributions of 11.28%, 9.62%, and 15.39%, respectively), while Hg, Cd, and Zn exhibit the greatest anti-bonding interactions with their nearest neighbor sulfur atoms (with percent anti-bonding contributions of 84.12%, 90.0%, and 85.21%, respectively); a visualization of overall trends is presented in Figure 6. Thus, the COHP analysis shows that there is greater correlation with the anti-bonding character of chalcogen in the adsorbate interaction and the adsorption energy. We attribute this to the fact that when the adatom gains or loses charge upon adsorption, that charge is exchanged with primarily the surrounding chalcogen atoms. For example, silver adsorbed onto WSe2, the chalcogen loses a total charge of 4.34 electrons compared to a selenium vacancy in WSe2 This finding also provides physical insights as to the selection of descriptors for random forest regression of adsorption energies in the work by Frey et al [43], where it was not clear why the mean number of p-valence electrons was ranked as the highest in importance. Here, our analysis reveals that the charge transfer from the anion atoms, whose valence electrons are primarily of p-character, is a large influence on the extent of electron filling of the anti-bonding orbitals, which correlates with the strength of adsorption interactions.
Figure 6: Comparison of COHP results describing the percentage of anti-bonding interactions of the adsorbate and sulfur (ab-ADS-S) and total anti-bonding interactions in MoS2. The anti-bonding interaction percentage between the adsorbate and sulfur is correlated with adsorption energy with an R2 value of 0.70, while the combined anti-bonding character has an R2 value of 0.57.
@@
We also conducted an analysis of the hybridization character of the adsorbates’ d-orbitals to evaluate whether differences in adsorption energy observed across different periods could be better understood. For adsorbates within the same group, despite their general similarity in electron configuration, the hybridization of the 3d, 4d, and 5d orbitals varies with the substrate. To quantify the amount of $d_{xy},d_{xz},d_{yz},d_{z^{2}}$, and $d_{x^{2}-y^{2}}$ orbital character, we integrate the projected density of states over a fixed energy range, from -2.0 eV to 0.5 eV relative to the Fermi level, to capture any mid-gap states introduced by the adsorbates; these data may be found in Figures S9 and S10 of the SI.
@@
Figure 7: The charge density difference between the adsorbed structure, the vS MoS2 structure, and the adsorbate plots of Nb, Y, and Zr, viewed along the b-axis and the c-axis. The yellow distribution corresponds to charge accumulation and the blue corresponds to charge depletion.
@@
Charge Density Difference (CDD) plots provide further insight into the nature of these interactions. For instance, in the case of Nb, Y, and Zr (shown in Figure 7), which all show a higher degree of $d_{yz}$ and $d_{xz}$ contributions. The CDD plots reveal regions of charge depletion near the adsorbate and charge accumulation near the substrate. This visual evidence supports the argument that these adsorbates have ionic bonding characteristics, where the spatial alignment of the $d_{yz}$ and $d_{xz}$ orbitals of the adsorbate enables greater interaction with the substrate. Overall, the variation in adsorption energies within the same group can be partially understood to arise from differences in the hybridization and spatial orientation of specific $d$-orbitals with the substrate orbitals.
III.3 Prediction of adsorption energies
@@
To explore the prediction of adsorption energies, we employ the Sure Independence Screening and Sparsifying Operator (SISSO) [72], a compressed sensing technique, to predict absolute adsorption energies using basic descriptors. In this context, dimensionality refers to the number of features or descriptors used to represent the data. High-dimensional datasets, where the number of potential descriptors is large, can be challenging because they increase the complexity of the model and the risk of overfitting in which the model becomes too tailored to the specific dataset and loses its ability to generalize to new data.
@@
SISSO addresses these challenges by first reducing the dimensionality of the feature space. It does this by discarding weakly correlated features through sure independence screening (SIS), focusing the analysis on the most relevant descriptors. This reduction is crucial for managing complexity and ensuring that the model remains both robust and generalizable. Next, the sparsifying operator (SO) identifies the optimal $n$-dimensional descriptor by selecting a minimal set of features that best predict the adsorption energies, ensuring that the model is both accurate and interpretable. Unlike many machine-learning techniques, SISSO can effectively perform non-linear regression on small datasets while preserving interpretability, making it ideal for complex datasets where transparency is essential. A list of the initial set of descriptors used in the SISSO analysis is provided in the SI.
@@
We leverage a dataset of readily available atomic properties for the substituents in each structure to construct the SISSO descriptor [72, 76]. In Section III.1, we demonstrated the significance of bonding and anti-bonding contributions to adsorption energies. This leads us to consider whether the d-band center calculation, which typically requires a detailed density of states (DOS) analysis, could be avoided. Since DOS calculations are a major computational bottleneck in predicting adsorption energy, identifying an alternative method or descriptor that bypasses this step could improve the computational efficiency of adsorption energy predictions.
To assess the performance of our SISSO-based approach, we compared the predicted adsorption energies to the values obtained via DFT calculations. Figure 8 shows the results of the best-performing compound descriptors using one- and three-dimensional descriptors. The SISSO analysis is performed for the specific case of MoS2 in Figure 8a as well as for the aggregated adsorption energies across four substrates (MoS2, MoSe2, WS2, and WSe2) in Figure 8b). When considering just the MoS2 substrate, the 1D SISSO descriptor yields an $R^{2}$ of 0.70, which improves slightly to 0.77 when increasing the dimensionality to 3D. A similar improvement is observed when comparing the 1D and 3D SISSO descriptors for all substrates considered. For the case of MoS2 alone, the 1D SISSO descriptor yields a mean absolute error (MAE) of 0.89 eV, a mean squared error (MSE) of 1.15 eV, and an R2 value of 0.71. The 3D SISSO descriptor for MoS2 improves these metrics, achieving an MAE of 0.76 eV, an MSE of 1.04 eV, and an R2 of 0.77.
When considering all substrates in combination (MoS2, MoSe2, WS2, and WSe2), the 1D SISSO descriptor results in an MAE of 0.97 eV, an MSE of 1.44 eV, and an R2 of 0.64. In comparison, the 3D SISSO descriptor for the same set achieves an improved MAE of 0.86 eV, an MSE of 1.31 eV, and an R2 of 0.70.
Figure 8: Prediction of adsorption energies from best-performing 1D and 3D SISSO descriptors for (a) MoS2 where reference DFT values are in black and SISSO values are in colored, and (b) MoS2, MoSe2, WS2, and WSe2 substrates.
Compared to other computational studies on the high-throughput calculation of defect and adsorption energies, our error metrics provide some insight in degree of error introduced when considering a minimal set of simple descriptors compared to more complex and computationally expensive non-linear regression. For instance, a study on defect formation energies in the zinc blende materials class, using random forest regression, Gaussian process regression, and neural network regression, achieved mean squared errors (MSEs) ranging from 0.94 to 1.05 eV for formation energies up to approximately 11 eV [77]. Another data-driven study on the defect formation energy of oxygen vacancies in metal oxides reported MSEs ranging between 0.21 and 0.45 eV using a suite of machine-learning regression models [78]. The approach with SISSO demonstrates promising accuracy with using only primary atomic properties—such as atomic radii and electronegativities—without the need for computationally expensive DOS calculations. This comparison highlights the efficiency of our method, such as in high-throughput studies where balancing accuracy and computational cost is crucial. It also highlights the challenges in predicting defect energetics to within chemical accuracy and emphasizes the value of a physics-informed descriptor.
@@
Finally, we comment on the anticipated transport characteristics of the adsorbed metal adatoms, a pre-requisite for such point defects to be used in resistive switching contexts. Adsorption of gold onto sulfur vacancies in MoS2 is known to lead to up to around ten orders of magnitude of change in the normalized conductivity, including the out-of-plane component, depending on the strain present [34]. The change in conductivity is accompanied by a partially-filled defect level with predominantly Au $d$-character that imbues a degree of metallic character to the substrate. Hence, the change in conductivity can, to first order, be explained by the increase in density of states near the Fermi level [34]. When inspecting the DOS of the adsorbates studied here, we observe that 76% of the configurations tested are expected to show enhanced metallic-like behavior with the adatom adsorbed (a list of the elements may be found in Table S1 of the SI). Further investigation of the transport properties is merited and the topic of future study.
III.4 Physical model for resistive switching based on point defects
Building on the computed adsorption energies, we establish a connection between these energies and the switching energy in vertically-stacked memristors composed of monolayer TMDs. Our model of the switching energy is based on the dissociation-diffusion-adsorption (DDA) model [33, 14] (Figure S1). In the DDA model, an adatom from the electrode dissociates and then potentially diffuses across the surface before adsorbing to a pre-existing anion vacancy (Figure S1). This process involves an energy of dissociation, energy of diffusion, and energy of adsorption. This mechanism is supported through previous measurements of single-defect memristor in MoS2 based on local scanning tunneling microscopy images [46]. The energy of dissociation is determined by the chemical identity of the electrode material, and thus is assumed to be constant for a given adatom species. A full study of diffusion is outside the scope of this paper, but is the topic of on-going study. Thus, the main quantity of interest is the adsorption energy.

### v2

@@ 2.2 Rationalizing trends in adsorption energy
Physically, the strength of interaction between the adsorbate and substrate is related to the filling of the anti-bonding states between the re-normalized valence states of the substrate and adsorbate upon adsorption. Because anti-bonding states generally lie energetically above the electronic band edge, the center of the d-states is an indicator of the bond strength between adsorbate and substrate. The center of the d-band is often referenced to the Fermi level, which determines the energy level of half occupation and thus delineates the filled and unfilled electronic states. Thus, a higher in energy d-band center relative to the Fermi level is indicative of a stronger adsorption interaction, as fewer anti-bonding states are occupied. Hence, the d-band model provides a basis for explaining differences and trends in adsorption energetics.
Traditionally, this model focuses on the position of the d-band center of the transition metal substrate relative to the Fermi level as the typical adsorbates are organic chemical intermediates involved in catalytic reactions. By contrast, the adsorbates in this study are transition metal adatoms on select substrates (MoS2, MoSe2, WS2, WSe2) that show minimal difference in their structural relaxations (Figure 8). In this study, we apply the d-band center model by focusing on the d-band centers of the transition metal adsorbates instead of the TMD substrates. This adaptation of the d-band center model leverages the underlying principle that the orbital filling and the position of the d-band center describe differences in adsorption energies. This dependence on the electronic structure is illustrated in Equations 2 and 3. Studies such as Qiao et al[49] have also investigated hydrogen adsorption on single-atom-anchored 2D basal planes, finding a correlation between the $d_{z^{2}}$-band center and adsorption strength. These observations collectively suggest that similar principles may govern interactions involving transition metal adatoms on low-dimensional materials.
@@
As observed from Figure 3 for MoS2, the early transition metals, such as Hf, Zr, and Ti, have d-band centers shifted closest to the Fermi level when compared to other transition metals. A similar observation can be made for MoSe2, WS2, and WSe2 (see Supplementary Figures 8, 9 , and 10). According to the d-band model, this shift of the d-bands closer to the Fermi level correctly anticipates a stronger adsorption energy, as the anti-bonding states would be minimally occupied. In contrast, the late transition metals, such as Cd, Hg, Zn, Au, Cu, and Ag, have d-band centers shifted farther from the Fermi level, indicating lower energy levels for the anti-bonding orbitals, leading to weaker adsorption energy. Consequently, the d-band centers for the re-normalized electronic states of the middle transition metals are shifted by an intermediate amount, resulting in intermediate adsorption energies and strengths. By comparing the shifts of the d-band centers of different adsorbates on the surfaces of TMDs, we can directly correlate the electronic structure of the adsorbates with their adsorption energies. This approach provides a clear understanding of the trends observed across different transition metal adsorbates, highlighting the utility of the d-band center model in rationalizing trends in adsorption properties in resistive switching contexts [50].
Interestingly, we observe that by separating the transition metal adsorbates into 3d, 4d, and 5d metals, there is a more robust correlation between the computed adsorption energy and the d-band center (Figure 3). We hypothesize that this difference in correlation based on period stems from the increasing spatial extents of the 3d, 4d, and 5d orbitals. An examination of the hybridization characteristics of the adsorbate reveals no discernible trends (see Supplementary Figures 11 and 12). Deviations from the linear d-band model are known, such as when multiple transition metal species are involved as in skin alloys [51]. The multiplicity in spin state of transition metals has also been shown to break linearity in adsorption energetics [52], particularly for the spatially-localized, highly-correlated 3d metals where higher-accuracy wavefunction-based methods are often required to capture and resolve different spin states [52, 53]. Supplementary Figures 8, 9, and 10 provide equivalent data for MoSe2, WS2, and WSe2.
@@
We next turn to crystal orbital Hamilton population (COHP) analysis. COHP analysis provides insights into the bonding and anti-bonding character in the interactions between adsorbates and substrates, helping to further understand the nature of adsorption. By decomposing the electronic density of states (DOS) into its bonding and anti-bonding components as a function of energy, we recover trends in line with the d-band center model. Notably, our COHP analysis reveals that anti-bonding interactions involving sulfur atoms play a role in the adsorption energy trends. Figure 4 shows examples for how the degree of anti-bonding character is computed based on COHP analysis. From this analysis, we observe that an increasing contribution of anti-bonding interactions from molybdenum and sulfur leads to a destabilization of adsorption energies.
We used our COHP analysis to decompose the overall bonding and anti-bonding interactions by species. Figure 5 shows that percent sulfur anti-bonding character and adsorption energy are correlated ($R^{2}\sim$ 0.70). This analysis is in line with the d-band model, in which the percent anti-bonding character of states below the Fermi level correlates with inversely with the adsorption strength, and further reveals the species contributions. For additional examples, the COHP analysis shows that Hf, Zr, and Ti have the lowest anti-bonding interactions with their nearest neighbor sulfur atoms (with percent anti-bonding contributions of 11.28%, 9.62%, and 15.39%, respectively), corresponding to generally stronger adsorption. Hg, Cd, and Zn exhibit the greatest anti-bonding interactions with their nearest neighbor sulfur atoms (with percent anti-bonding contributions of 84.12%, 90.0%, and 85.21%, respectively), corresponding to generally weaker adsorption.
These findings highlight that adsorption strength is governed by multiple factors, including charge redistribution and antibonding interactions. A comparison of the correlation between adsorption energy with Bader charge, percent Mo anti-bonding character, and percent S anti-bonding character is provided in Supplementary Figure 13. This finding also provides physical insights as to the selection of descriptors for random forest regression of adsorption energies in the work by Frey et al. [43], where it was not clear why the mean number of p-valence electrons was ranked as the highest in importance. Here, our analysis reveals that the charge loss or gain from the chalcogen atoms, whose valence electrons are primarily of p-character, strongly influences the filling of anti-bonding orbitals, correlating with adsorption strength.
@@
We also conducted an analysis of the hybridization character of the adsorbates’ d-orbitals to evaluate whether differences in adsorption energy observed across different periods could be better understood. For adsorbates within the same group, despite their general similarity in electron configuration, the hybridization of the 3d, 4d, and 5d orbitals varies with the substrate. To quantify the amount of $d_{xy},d_{xz},d_{yz},d_{z^{2}}$, and $d_{x^{2}-y^{2}}$ orbital character, we integrate the projected density of states over a fixed energy range, from -2.0 eV to 0.5 eV relative to the Fermi level, to capture any mid-gap states introduced by the adsorbates; these data may be found in Supplementary Figures 11 and 12. This range was chosen to focus on the energy window most relevant to resistive switching and charge transport, where adsorbate-induced hybridization with defect sites is expected to be strongest. It includes the entire band gap and portions of the conduction and valence bands, ensuring contributions from both filled and empty hybrid states are accounted for. While deeper integration windows would include broader bonding states, we found that the relative orbital contributions remain largely consistent across different windows for most adsorbates.
@@
Charge density difference (CDD) plots provide further insight into the nature of these interactions. For instance, Nb, Y, and Zr are shown in Figure 6, which all show a higher degree of $d_{yz}$ and $d_{xz}$ contributions. The CDD plots reveal regions of charge depletion near the adsorbate and charge accumulation near the substrate. This visual evidence supports the argument that these adsorbates have ionic bonding characteristics, where the spatial alignment of the $d_{yz}$ and $d_{xz}$ orbitals of the adsorbate enables greater interaction with the substrate.
Overall, the variation in adsorption energies within the same group can be partially understood to arise from differences in the hybridization and spatial orientation of specific $d$-orbitals with the substrate orbitals.
2.3 Prediction of trends in adsorption energies
@@
To explore the prediction of adsorption energies, we employ the Sure Independence Screening and Sparsifying Operator (SISSO) [54], a compressed sensing technique to predict adsorption energies using basic descriptors. In this context, dimensionality refers to the number of features or descriptors used to represent the data. High-dimensional datasets, where the number of potential descriptors is large, can be challenging because they increase the complexity of the model; there is also a risk of overfitting, in which the model becomes too tailored to the specific dataset and loses its ability to generalize to new data.
@@
SISSO addresses these challenges by first reducing the dimensionality of the feature space discarding weakly correlated features through sure independence screening (SIS), focusing the analysis on the most relevant descriptors. This reduction is crucial for managing complexity and ensuring that the model remains both robust and generalizable. Next, the sparsifying operator (SO) identifies the optimal $n$-dimensional descriptor by selecting a minimal set of features that best predict the adsorption energies, ensuring that the model is both accurate and interpretable. Unlike many machine-learning techniques, SISSO can effectively perform non-linear regression on small datasets while preserving interpretability, making it ideal for complex datasets where transparency is essential. A list of the initial set of descriptors used in the SISSO analysis is provided in the SI.
@@
We leverage a dataset of readily available atomic properties for the substituents in each structure to construct the SISSO descriptor [54, 55]. In Section 2.1, we demonstrated the significance of bonding and anti-bonding contributions and adsorbate charge loss or gain to adsorption energies. This leads us to consider whether the d-band center calculation, which typically requires a detailed density of states (DOS) analysis, could be avoided with proxy descriptors. Identifying an alternative method or descriptor that bypasses this step could improve the computational efficiency of adsorption energy predictions.
To assess the performance of our SISSO-based approach, we employed a leave-one-out (LOO) validation strategy. In this approach, one TMD substrate was entirely excluded from the training set and used as an independent test case, repeating the process for (MoS2, MoSe2, WS2, and WSe2. The complexity of our SISSO models is governed by user-defined parameters, such as descriptor dimensionality and the allowed complexity in combining features (see Supplementary Section 3).
Figure 7 presents the predicted versus reference DFT adsorption energies for the best-performing 5D SISSO model across different LOO cases. For instance, when trained on MoSe2, WS2, and WSe2, the model was tested on MoS2 (Figure 7b and c). Supplementary Table 2 quantifies model performance across all LOO scenarios, summarizing training and test set R2, mean squared error (MSE), and mean absolute error (MAE). The SISSO model maintains test set R2 values ranging from 0.789 to 0.892 across all LOO cases. We note that even 1D SISSO models can achieve R2 = 0.64 (Supplementary Figure 18). Physical quantities, including electronegativity ($X$), ionization energy ($IE$), work function ($WF$), atomic radius ($R_{c}$), and valency, were consistently retained in all models.
SISSO models that include the d-band center and structural information (e.g., average distance between adsorbate and neighboring molybdenum atoms) were also evaluated (see Supplementary Figures 21 and 22). These models provide further improvements in R2 up to 0.91. Increasing the dimensionality beyond 5D resulted in diminishing returns. Simpler models that explicitly account for some known limitations of the d-band center model also improve the descriptive nature of the d-band center model. Supplementary Figures 19 and 20 show that including the covalent radius information about the atomic species in each system, structural information in the form of average distance between the adsorbate and substrate, and filling information (i.e., number of valence electrons) improves the correlation and descriptive nature of the d-band center model. In both 2D and 3D SISSO models, utilizing the d-band center, structural information, filling information, and categorical descriptors based on periodic blocks, we achieve R2 values of 0.80 and 0.84 (Supplementary Figures 19 and 20), respectively). Overall, the agreement between SISSO-predicted and DFT reference values (Figure 7) and the relative consistency in R2, MAE, and MSE across LOO test cases (Supplementary Table 2) suggest that atom-based descriptors can be used to predict trends in adsorption energetics in TMDs.
Compared to other computational studies on high-throughput defect and adsorption energy calculations, our error metrics provide insight into the degree of error introduced when using a minimal set of simple descriptors, such as atomic radii and electronegativities, in contrast to more complex and computationally expensive non-linear regression methods. For example, Mannodi-Kanakkithodi and colleagues [56] reported mean squared errors (MSEs) of 0.94–1.05 eV for defect formation energies up to 11 eV in zinc blende materials using advanced machine learning models like random forests and neural networks. Similarly, Wan et al[57] observed MSEs of 0.21–0.45 eV for oxygen vacancy formation energies in metal oxides using various multiple linear and machine-learning regression models. These studies contain regression models that require large amounts of data and have limited interpretability. Frey et al[43] used ensemble random forest regression to achieve R2= 0.74 and MAE = 0.67 in adsorption energies of metal adatoms in TMD monolayers. The top two out of ten reported descriptors included the mean number of $p$-valence electrons and the chemical potential. The chemical potential is already a term in the calculation of the adsorption energy, and reveals little of the adsorbate-absorbent interactions. Our previous COHP analysis reveals that the interactions of the adsorbate with surrounding chalcogen atoms may partially explain the descriptor of mean $p$-valence electrons.
In comparison, the SISSO approach offers reasonable accuracy for simpler models with smaller datasets. Here, we choose descriptors that directly link to basic chemical parameters (e.g., atomic radii, electronegativity, ionization energies). We show that models containing one descriptor (e.g., $d$-band center) to models based on three to five descriptors can explain a majority of the trends in the adsorption energy (R2 = 0.89 to 0.91). The simplicity of our SISSO models leads to a tradeoff of slightly higher deviations from reference data, as seen in our adsorption energy calculations (Figure 7a). Nevertheless, these models serve as a valuable starting point, balancing computational efficiency with model simplicity and reasonable predictive accuracy, particularly in high-throughput studies where efficiency is critical or in resource-limited computational studies.
@@
Finally, we comment on the anticipated transport characteristics of the adsorbed metal adatoms, a pre-requisite for such point defects to be used in resistive switching contexts. Adsorption of gold onto sulfur vacancies in MoS2 is known to lead to up to around ten orders of magnitude of change in the normalized conductivity, including the out-of-plane component, depending on the strain present [34, 14, 46]. The change in conductivity is accompanied by a partially-filled defect level with predominantly Au $d$-character that imbues a degree of metallic character to the substrate. Notably, calculations from Anvari and Wang [34] predicted a ratio of around 10 orders of magnitude difference between the low- and high-resistive states, which compares comparably with the experimentally observed ratio of 686-86 - 8 orders of magnitude difference [46]. Hence, the change in conductivity can, to first order, be explained by the increase in density of states near the Fermi level [34]. When inspecting the DOS of the adsorbates studied here, we observe that 76% of the configurations tested are expected to show enhanced metallic-like behavior with the adatom adsorbed (a list of the elements may be found in Supplementary Table 1). Further investigation of the transport properties is merited and the topic of future study.
2.4 Physical model for resistive switching based on point defects
Building on the computed adsorption energies, we establish a connection between these energies and the switching energy in vertically-stacked memristors composed of monolayer TMDs. Our model of the switching energy is based on the dissociation-diffusion-adsorption (DDA) model [33, 14] (Supplementary Figure 1).
In the DDA model, an adatom from the electrode dissociates diffuses across the surface before adsorbing to a pre-existing anion vacancy (Supplementary Figure 1). This process involves three key energy components: dissociation energy, diffusion energy, and adsorption energy.
We observe a correlation (R2 = 0.96) between the calculated adsorption energies of various metal adatoms on monolayer TMDs and the experimentally observed set power in corresponding memristor devices (Supplementary Figure 2). Specifically, stronger (more negative) adsorption energies are associated with reduced set power. This correlation suggests the adsorption energy plays a role in the switching energetics of the device. The energy of dissociation is determined by the chemical identity of the electrode material and thus is assumed to be constant for a given adatom species. An additional analysis comparing calculated adsorption energies with the cohesive (bulk) energies of the corresponding metals is shown in Supplementary Figure 3. This comparison reveals a strong negative correlation ($R^{2}\approx 0.80$), indicating that metals with higher cohesive energies typically exhibit stronger adsorption energies. A full study of diffusion is outside the scope of this paper, but is the topic of ongoing study. Thus, our focus from an energetics point of view is the adsorption energy.
The DDA model is supported by previous measurements of single-defect memristors in MoS2 based on local scanning tunneling microscopy images [46]. In these studies, gold electrodes were observed to dissociate into adatoms, which then interact with the MoS2 surface, modulating the resistive switching state. Different electrode materials can lead to distinct switching mechanisms. For instance, devices utilizing silver electrodes often exhibit filamentary switching behaviors, where the formation and dissolution of conductive filaments play a crucial role in the resistive switching process [58, 59, 60]. A more systematic comparison across electrode materials is required, for which this study provides initial materials-based insights into expected trends.